OpenStudy (anonymous):

Solve for N 9/8 - (5n)/(4n+8) = 2/(n+2)

OpenStudy (anonymous):

n =2

OpenStudy (anonymous):

9/8=2/(n+2)+5n/(4(n+2) 9/8=8/(4(n+2))+5n/(4(n+2)) 9/8=(8+5n)/(4n+8) 9(4n+8)=8(8+5n) 36n+72=64+40n 8=4n n=2

OpenStudy (anonymous):

note:the n+2 factor gets cancelled out.